A tire is rotating 900times per min. Through how many degrees does a point on the edge of a tire move in 1/3 seconds?

Found 2 solutions by lwsshak3, Alan3354:
Answer by lwsshak3(6463)

A tire is rotating 900times per min. Through how many degrees does a point on the edge of a tire move in 1/3 seconds?
**
900rev/min*360º/rev*min/60 sec=900*360º/60 sec=5400º/sec
Degrees a point on the edge of the tire move in 1/3 seconds=1/3*5400º/sec=1800º

Answer by Alan3354(30968)

A tire is rotating 900times per min. Through how many degrees does a point on the edge of a tire move in 1/3 seconds?
-----------------
900 rev/min = 15 rev/sec
--> 5 rev in 1/3 second
= 1800 degrees
=================
For all points on the tire (except the center), not just a point on the edge.